# Relations between average clustering coefficient and another centralities in graphs

### v1

@@ Abstract
    $c_{i}=c(i)=\frac{\text{number of edges in }N(i)}{\text{maximum possible numberof edges in }N(i)}=\frac{2\|E(N(i)))\|}{d_{i}(d_{i}-1)}$.
@@
    $C_{WS}(G)=\frac{1}{n}\sum\limits_{i\in V(G)}c_{i}=\frac{1}{n}\sum\limits_{i\in V(G)}\frac{2\|E(N(i)))\|}{d_{i}(d_{i}-1)}=\frac{1}{n}\sum\limits_{i\in V(G)}\frac{\sum\limits_{j,k\in V(G)}a_{ij}a_{jk}a_{ki}}{d_{i}(d_{i}-1)}$.
@@
First let’s prove a lemma about a relation between average shortest path length between vertices in the neighbourhood of $i$ and local clustering coefficient of this vertex.
@@
$=\frac{1}{d_{i}(d_{i}-1)}(2\|E(N(i))\|+\sum\limits_{(s,i),(i,t)\in E(G),(s,t)\notin E(G)}dist(s,t))=$
$=\frac{1}{d_{i}(d_{i}-1)}(2\|E(N(i))\|+2(d_{i}(d_{i}-1)-2\|E(N(i))\|))=2-c_{i}.$
@@
Let’s prove theorem about a relation between average clustering coefficient and betweenness centrality.
@@
$BC(i,N(i))=\sum\limits_{j,k\in N(i),\;(j,k)\notin E(N(i))}\frac{1}{\sigma_{jk}}\leq\sum\limits_{j,k\in N(i),\;(j,k)\notin E(N(i))}1=d_{i}(d_{i}-1)-2\|E(N(i))\|,$
@@
Note that for $diam(G)=2$ holds an equality. ∎
@@
Now let’s prove theorem about a relation between average clustering coefficient and radiality.
@@
$C_{WS}(G)=\frac{1}{n}\sum\limits_{i\in V(G)}(\frac{1}{d_{i}}\sum\limits_{v\in N(i)}\overline{Rad}(v)-1)+\frac{\#\{N(i)\text{ which are complete graphs}\}}{n}.$
@@
$\frac{1}{d_{i}}\sum\limits_{v\in N(i)}\overline{Rad}(v)=diam(N^{\prime}(i))+1-L(N(i))=diam(N^{\prime}(i))-1+c_{i}=c_{i}+1-\chi_{K_{d_{i}}}(N^{\prime}(i)),$
where $\chi_{K_{d_{i}}}(N^{\prime}(i))=\begin{cases}1&\text{if $N^{\prime}(i)=K_{d_{i}}$}\\ 0&\text{otherwise}\end{cases}$. Averaging by $i$ ends the proof. ∎
@@
Let’s $\forall i,j\in V(G),\;i\leq j$ hold $d_{i}\leq d_{j}\Rightarrow c_{i}\leq c_{j}$, then
@@
The equality holds when $\forall i,j\in V(G):d_{i}=d_{j}$, that is for a graphs in which all vertices degrees are equal. ∎
@@
Let’s $\forall i,j\in V(G),\;i\leq j$ hold $d_{i}\leq d_{j}\Rightarrow c_{i}\geq c_{j}$, then

### v2

@@ Abstract
    $c_{i}=c(i)=\frac{\text{number of edges in }N(i)}{\text{maximum possible numberof edges in }N(i)}=\frac{2|E(N(i))|}{d_{i}(d_{i}-1)}$.
@@
    $C_{WS}(G)=\frac{1}{n}\sum\limits_{i\in V(G)}c_{i}=\frac{1}{n}\sum\limits_{i\in V(G)}\frac{2|E(N(i))|}{d_{i}(d_{i}-1)}=\frac{1}{n}\sum\limits_{i\in V(G)}\frac{\sum\limits_{j,k\in V(G)}a_{ij}a_{jk}a_{ki}}{d_{i}(d_{i}-1)}$.
@@
Let’s define in the similar way as for local efficiency the local average shortest path length, local betweenness, local stress and local radiality centralities.
Let’s
  1. 1.
    $L\bigl{(}N(i)\bigr{)}=\frac{1}{d_{i}(d_{i}-1)}\sum\limits_{v,w\in N(i)}\operatorname{dist}(v,w)$ be the average shortest path length for vertices of the neighborhood $N(i)$, where the average shortest path length is defined in the whole graph $G$,
  2. 2.
    $BC(i,N(i))=\sum\limits_{s,t\in N(i),\;s\neq t}\frac{\sigma_{st}(i)}{\sigma_{st}}$ be the betweenness centrality for vertices of the neighborhood $N(i)$, where shortest paths are defined in the whole graph $G$, and $BC_{loc}(G)=\frac{1}{n}\sum\limits_{i\in V(G)}\frac{BC(i,N(i))}{d_{i}(d_{i}-1)}$the local betweenness centrality,
  3. 3.
    $\operatorname{Rad}(v,N(i))=\frac{\sum\limits_{t\in N(i),t\neq v}\big{(}\operatorname{diam}(N(i))+1-\operatorname{dist}(v,t)\big{)}}{d_{i}-1}$ be the radiality centrality for vertices of the neighborhood $N(i)$, where shortest paths and the diameter are defined in the whole graph $G$, and $\operatorname{Rad}_{loc}(G)=\frac{1}{n}\sum\limits_{i\in V(G)}\frac{\sum\limits_{v\in N(i)}\operatorname{Rad}(v,N(i))}{d_{i}}$the local radiality centrality.
Let’s prove a lemma about a relation between average shortest path length between vertices in the neighborhood of $i$ and local clustering coefficient of this vertex.
@@
$=\frac{1}{d_{i}(d_{i}-1)}(2|E(N(i))|+\sum\limits_{(s,i),(i,t)\in E(G),(s,t)\notin E(G)}dist(s,t))=$
$=\frac{1}{d_{i}(d_{i}-1)}(2|E(N(i))|+2(d_{i}(d_{i}-1)-2|E(N(i))|))=2-c_{i}.$
@@
Let’s prove theorem about a relation between average clustering coefficient and local betweenness centrality.
@@
$BC(i,N(i))=\sum\limits_{j,k\in N(i),\;(j,k)\notin E(N(i))}\frac{1}{\sigma_{jk}}\leq\sum\limits_{j,k\in N(i),\;(j,k)\notin E(N(i))}1=d_{i}(d_{i}-1)-2|E(N(i))|,$
@@
Note that the equality holds, if there exists a unique shortest path between any two vertices of the neighbourhood $N(i)$. This holds if and only if there are no average shortest paths of the length 2 in $N(i)$, and thus $N(i)$ should be the disjoint union of several complete graphs for any vertex $i$. ∎
@@
Let’s prove theorem about a relation between average clustering coefficient and local radiality.
@@
$C_{WS}(G)=\operatorname{Rad}_{loc}(G)-1+\frac{\#\{N(i)\text{ which are complete graphs}\}}{n}.$
@@
$\frac{1}{d_{i}}\sum\limits_{v\in N(i)}\operatorname{Rad}(v,N(i))=\operatorname{diam}(N(i))+1-L(N(i))=\operatorname{diam}(N(i))-1+c_{i}=c_{i}+1-\chi_{K_{d_{i}}}(N(i)),$
where $\chi_{K_{d_{i}}}(N(i))=\begin{cases}1&\text{if $N(i)=K_{d_{i}}$}\\ 0&\text{otherwise}\end{cases}$. Averaging by $i$ ends the proof. ∎
@@
Let’s $\forall i,j\in V(G)$ hold if $d_{i}\leq d_{j}$ then $c_{i}\geq c_{j}$, then
@@
The equality holds when $\forall i,j\in V(G):d_{i}=d_{j}$ (i.e. for regular graphs) or when $\forall i,j\in V(G):c_{i}=c_{j}$. ∎
@@
Let’s $\forall i,j\in V(G)$ hold if $d_{i}\leq d_{j}$, then $c_{i}\geq c_{j}$, then
@@
Corollary 3.
For simple regular graphs $G$:
$C_{WS}(G)=C(G).$
For most well-known graphs $C_{WS}(G)\geq C(G)$, but it is not very hard to come up with an example then $C_{WS}(G)<C(G).$ Consider the complete graph $K_{n}$ and glue to each vertex cycle of the length 4. Then, the vertices of the complete graph have by symmetry the same $c_{i}>0$ and new added vertices have $c_{i}=0$. Thus, by theorem 6 for new graph $C_{WS}(G)<C(G).$